OpenStudy (anonymous):

Sam bought 2 1/4 pounds of Swiss cheese and 1 1/2 pounds of cheddar at the deli. How many pounds of cheese did Sam buy altogether?

OpenStudy (hoblos):

2 1/4 + 1 1/2 = 9/4 + 3/2 = 9/4 + 6/4 = 15/4 = 3 3/4

OpenStudy (anonymous):

thanks alot :)

OpenStudy (hoblos):

my pleasure XD

OpenStudy (anonymous):

;)